OpenStudy (anonymous):

WILL GIVE MEDAL!!! Which percent is equivalent to 3 2/5 ? A. 3.4% B. 32.5% C. 325% D. 340%

OpenStudy (anonymous):

@sammixboo @cookiemonster2412

OpenStudy (danjs):

turn to improper fraction 17/5 . multiply by 20/20 , to get a #/100

OpenStudy (danjs):

that # is the percent

OpenStudy (anonymous):

??

OpenStudy (danjs):

\[\frac{ 17 }{ 5 }*\frac{ 20 }{ 20 }=\frac{ ? }{ 100 }\]

OpenStudy (anonymous):

@Directrix @grrrxxiii @kohai @Forcequest @whpalmer4 @Abhisar @gracekygirl @dan815

OpenStudy (anonymous):

im to dumb to understand @DanJS

OpenStudy (anonymous):

5 means the number of times you will split the whole. a whole is equal to 100% so you will divide 100 by 5.

OpenStudy (anonymous):

Im late but correct answer is A: 3.4%

OpenStudy (anonymous):

My bad I meant D 340%